# Lightweight Deep Learning-Based Channel Estimation for RIS-Aided Extremely Large-Scale MIMO Systems on Resource-Limited Edge Devices

### v1

@@ I-B Contributions
This is the first work to address the challenges of cascaded channel estimation for RIS-assisted XL-MIMO systems. More specifically, the key contributions of our work are summarized as follows:
@@
  • We propose a UNet 3+-based architecture, customized with a dedicated denoising module to effectively extract noise from the channel coefficients and enable efficient cascaded channel estimation. The proposed architecture demonstrates reliable performance even when trained using limited data, maintaining high accuracy with small patch sizes. This highlights its robustness, computational efficiency, and adaptability for deployment on low-resource edge devices.
@@
Notation: $T$ and $H$ represent the transpose and Hermitian transpose of a matrix, while $\mathbb{R}$ and $\mathbb{C}$ denote the sets of real and complex numbers, respectively. The function $\text{diag}(\cdot)$ constructs a diagonal matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm, and $e$ indicates the exponential function. ${\mathbf{I}_{M}}$ represents the identity matrix of size $M\times M$. The operator $(\cdot)^{-1}$ denotes the matrix inverse, while $\mathbb{E}(\cdot)$ indicates the statistical expectation. The symbol $\otimes$ is used to represent the Kronecker product.
@@ II System Model
We consider a real-world scenario where a signal propagates through a three-dimensional (3D) space, with a local spherical coordinate system defined such that $\phi$ represents the azimuth angle and $\psi$ denotes the elevation angle [18]. The reflecting and antenna elements are arranged in a two-dimensional grid within the YZ plane. Each element has $d_{H}$ and $d_{V}$ corresponding to the horizontal and vertical spacing, respectively. To systematically index the elements, they are ordered row by row and denoted as $n\in[1,N]$ RIS elements and $m\in[1,M]$ antenna elements. The position of the $n$-th element relative to the origin is:
@@
[…]
@@
$\boldsymbol{\omega}(\phi,\psi)=\frac{2\pi}{\lambda}\begin{bmatrix}\cos(\psi)\cos(\phi)\\ \cos(\psi)\sin(\phi)\\ \sin(\psi)\end{bmatrix},$ (4)
@@
$\mathbf{f}_{k}=\sqrt{\frac{\eta_{r}}{\eta_{r}+1}}\mathbf{f}^{{LoS}}_{k}+\sqrt{\frac{1}{\eta_{r}+1}}\mathbf{f}^{{NLoS}}_{k},$ (5)
@@
where $\alpha_{n}$ and $\gamma_{n}$ are the indices of reflecting elements in the horizontal and vertical directions, respectively. We consider the Non-LoS (NLoS) channels as quasi-static Rayleigh fading, which remains constant for the coherence interval. The NLoS channel between the RIS and the user is modeled using correlated Rayleigh fading [25]. The NLoS channel from the RIS to the user is given by:
@@ II-A Channel Estimation Framework
The system’s performance is heavily dependent on the availability of precise channel information. We adopted a Time Division Duplex (TDD) protocol, which exploits channel reciprocity to estimate downlink CSI from uplink CSI. The TDD protocol divides communication into time slots, known as coherence intervals, during which the channel response remains approximately constant [25]. We divide the channel estimation framework into two phases: direct channel estimation and cascaded channel estimation. In the first phase, the RIS is turned off, so no reflections occur. The direct channel between the user and the BS is estimated during this stage, as shown in Fig. 2.
@@
Moreover, in the second phase, we estimate the cascaded channel between the user, the RIS, and the BS. Estimating the cascaded channel requires $L$ distinct phase shift matrices, where $L\geq N$, to facilitate accurate channel estimation. For each phase shift matrix, each user is assigned an orthogonal pilot sequence $\mathbf{x}_{k}$ of length $u$, where $u\geq K$, ensuring the pilot’s orthogonality for accurate estimation. For the $k$-th user during the $l$-th subframe, the pilot signal received at the BS, denoted as $\mathbf{y}_{k}\in\mathbb{C}^{M\times 1}$, is given by the following expression:
@@
$\mathbf{y}_{k}=\frac{1}{Pu}(\mathbf{b}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{H}\text{diag}(\mathbf{\Theta}_{l})\mathbf{f}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{n}_{k}\mathbf{x}_{k}),$ (15)
@@
$\mathbf{y}_{k}=\mathbf{b}_{k}+\mathbf{H}\text{\text{diag}}(\mathbf{\Theta}_{l})\mathbf{f}_{k}+\mathbf{v}_{k},$ (16)
@@
where $\mathbf{G}_{k}=\ \mathbf{H}\,\text{diag}(\mathbf{f}_{k})\in\mathcal{C}^{M\times N}$ represents the cascaded channel and $\mathbf{V}_{k}=[\mathbf{v}_{k,1},\mathbf{v}_{k,2},\ldots,\mathbf{v}_{k,L}]$. According to [14], designing the optimal phase shift matrix $\mathbf{S}$ using a discrete Fourier transform maximizes the received signal power at the BS, which can be expressed as follows:
@@
where $W=e^{j\frac{2\pi}{L}}$ and $\mathbf{S}\mathbf{S}^{H}=L\mathbf{I}_{N}$. Cascaded channel estimation involves recovering $\mathbf{G_{k}}$ from the received pilots. The simplest approach is the LS method [26], which performs effectively when the channel is assumed to be static. The LS representation of the cascaded channel can be expressed as $\mathbf{\tilde{G}}_{k}^{LS}=\mathbf{Y}_{k}\mathbf{S}^{H}(\mathbf{S}\mathbf{S}^{H})^{-1}$. However, treating the channel as static fails to capture its spatial characteristics, especially when compared to data-driven methods. Moreover, under a Rayleigh fading model and assuming the availability of the statistical channel correlation matrix, the optimal estimator is the MMSE estimator. However, when the channel distribution is complex, as in the case of the cascaded channel in our scenario, the MMSE estimator becomes computationally expensive. To reduce this complexity, the MMSE estimator is often approximated by the LMMSE estimator. According to [27], the LMMSE estimator is given by:
@@
where $\mathbf{R}_{G_{k}}=\mathbb{E}(G_{k}^{H}G_{k})$ denotes the statistical channel correlation matrix. However, since $G_{k}$ corresponds to a cascaded channel, it does not follow the Rayleigh fading model. Consequently, the assumptions underlying the LMMSE estimator are violated, leading to a significant performance gap between the LMMSE estimation and the actual channel. To address this limitation, we propose a deep learning-based method that effectively learns the underlying channel structure and mitigates the performance gap introduced by the model mismatch in conventional methods.
@@ III Deep Learning Assisted Channel Estimation
Deep learning-assisted cascaded channel estimation has gained significant popularity in recent years [15, 11, 12, 13, 14, 16]. One of the key reasons for this growing interest is its extremely low inference time, which makes it well-suited for low-latency and URLLC applications. However, existing works primarily focus on small-scale scenarios with only a limited number of antennas and RIS elements. In this paper, we investigate the use of deep learning to estimate the channel in RIS-enabled XL-MIMO systems, which may involve hundreds of RIS elements and thousands of antennas. We leverage the spatial correlation inherent in the cascaded channel, resulting from the closely placed RIS elements and BS antennas. Building on this property, we propose a patch-based training framework that significantly reduces input dimensionality, enabling scalable training for large-scale systems on resource-constrained edge devices. This section outlines the proposed deep learning framework, while Sections IV and V offer in-depth discussions on spatial correlation and the patch-based training strategy.
@@ III-A Theoretical Background
Recent studies have highlighted that autoencoders can learn efficient, low-dimensional data representations by capturing essential patterns and structures while minimizing the reconstruction loss [22, 21, 23]. They also generalize very well, making them effective in modeling complex, high-dimensional distributions. Autoencoders also require less data to train because they learn compact latent representations that retain the most critical features of the input. This latent space not only enables efficient data compression and generalization, but also helps extracting meaningful features, which is particularly beneficial in our case. Despite their effectiveness, autoencoders may encounter reconstruction limitations, as compression into a bottleneck layer can lead to the loss of critical spatial features, hindering the accurate recovery of fine-grained structures. To address this challenge, the authors in [21] proposed the UNet architecture, which mitigates this issue through skip connections that directly transfer high-resolution features from the encoder to the decoder. This design enables UNet to preserve essential spatial correlations while effectively filtering out noise [23].
In our proposed framework, we leverage the UNet 3+ architecture, an advanced extension of UNet, designed to aggregate feature maps from multiple hierarchical levels. UNet 3+ improves upon the standard UNet by incorporating enhanced multi-scale feature aggregation, making it more effective for denoising tasks [24]. While traditional UNet relies on skip connections between corresponding encoder and decoder layers, UNet 3+ introduces full-scale skip connections, which aggregate feature maps from multiple levels instead of just one, as shown in Fig. 3. This results in richer contextual information, allowing the model to capture both fine-grained local details and broader structural patterns more effectively. This enhanced feature aggregation significantly improves denoising performance by preserving critical spatial structures and more effectively suppressing noise, even in scenarios with limited training data. This makes UNet 3+ a better choice for denoising tasks in large-scale systems than the standard UNet.
Figure 3: The proposed UNet 3+ framework employs $l$ levels, each with a dedicated denoising block in the encoder. Filters double, and the resolution halves at each level. Fused encoder and decoder features serve as input to each decoder layer.
@@ III-B Proposed Model Architecture
We propose a UNet 3+ based architecture designed to effectively capture the underlying signal features in the dataset. The framework integrates a dedicated denoising module that employs a subtraction-based structure to enhance channel estimation by isolating the core signal components from additive noise. The architecture, illustrated in Fig. 3, adopts a multi-level encoder–decoder design to exploit spatial hierarchies to break the noise correlation, enabling more effective denoising and improved estimation accuracy. In the encoder, each level in the architecture consists of multi-layer Convolutional Neural Networks (CNNs) with distinct denoising blocks. The output of each higher-level block is downsampled by a factor of two and fed into the subsequent lower-level block, where the number of filters is doubled to capture increasingly complex features. We employ a MaxPooling layer with a (2,2) kernel in the encoder to downsample between levels. In the decoder, the output from lower-level blocks is upsampled to the current level, whereas the feature maps from the corresponding encoder blocks are downsampled to match that level. Additionally, the output from the same level in the encoder is concatenated with these other outputs, enabling effective multi-scale information fusion to preserve spatial structure. Moreover, we utilized bilinear interpolation for upsampling between levels in the network to smoothly increase the resolution of feature maps while maintaining spatial consistency. During the decoder phase, the filter size is reduced by half, which helps to refine the reconstructed output. Finally, an activation function is applied to the aggregated features before generating the final output, introducing nonlinearity and enhancing the model’s ability to learn complex mappings.
@@ III-C Denoising Block Structure
The CNN architecture is well-suited to extract underlying signal features from noisy data effectively. In [11], a subtraction-based mechanism was proposed to eliminate noise from the data, demonstrating effectiveness in denoising. However, this approach was limited to learning models in which the input and output have identical dimensions, a constraint that does not align with our proposed architecture. To address this limitation, the proposed denoising architecture introduces a $1\times 1$ convolutional layer at the beginning of each level in the encoder to align the input and output dimensions, thereby facilitating the application of the denoising mechanism. The denoising block architecture is composed of a CNN with $i$ layers, where each layer consists of a convolution operation, batch normalization, and a non-linear activation function, as illustrated in Fig. 3. The denoising block takes an input and applies a pointwise ($1\times 1$) convolution to align it with the network’s internal dimensions. The transformed input then passes through multiple convolutional layers. Finally, an element-wise subtraction is performed between the $l$-th level’s block input $\mathbf{I}_{l}$ and output $\tilde{\mathbf{z}}_{\theta}(\mathbf{I}_{l})$ to produce the denoised channel matrix, effectively preserving essential features. At the $l$-th level, it can be expressed as:
$\mathbf{Z}_{l}=\mathbf{I}_{l}-\tilde{\mathbf{z}}_{\theta}(\mathbf{I}_{l})$ (20)
where $\tilde{\mathbf{z}}_{\theta}$ denotes the convolutional neural network and $\theta$ represents its parameters. The primary challenge associated with cascaded channel estimation is the presence of additive white Gaussian noise (AWGN) [6]. The simplest estimator for this task is the LS method, as it does not require statistical information. However, channel estimation using LS is often inadequate, as it is highly sensitive to noise, particularly in low Signal-to-Noise Ratio (SNR) conditions. Although it may still provide a rough estimate, its performance significantly degrades in noisy environments. Therefore, we use the LS estimator as input to our model, as it allows for a coarse estimate of the cascaded channel [11]. This initial estimate enables our model to focus on refining the channel representation, rather than learning it from scratch, thereby improving convergence and overall performance. Moreover, to accommodate the complex-valued nature of the data, each sample is decomposed into its real and imaginary components and provided as separate input channels to facilitate effective feature extraction. Thus, the input sample at the first level is represented as $\mathbf{I}_{0}=\left[\text{Re}\left\{\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\right\},\ \text{Im}\left\{\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\right\}\right]$.

### v2

@@ I-B Contributions
To the best of our knowledge, this is the first work that employs deep learning to address the challenges of cascaded channel estimation in RIS-assisted XL-MIMO systems. Specifically, the key contributions of this work are as follows:
@@
  • We propose a multi-scale feature extraction and aggregation-based architecture, customized with a dedicated denoising module to effectively extract noise from the channel coefficients and enable efficient cascaded channel estimation. The proposed architecture demonstrates reliable performance even when trained using limited data, maintaining high accuracy with small-sized patches. This highlights its robustness, computational efficiency, and adaptability for deployment on low-resource edge devices.
@@
Notation: $T$ and $H$ represent the transpose and Hermitian transpose of a matrix, while $\mathbb{R}$ and $\mathbb{C}$ denote the sets of real and complex numbers, respectively. The function $\text{diag}(\cdot)$ constructs a diagonal matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm, and $e$ indicates the exponential function. $\mathbf{I}_{M}$} represents the identity matrix of size $M\times M$. The operator $(\cdot)^{-1}$ denotes the matrix inverse, while $\mathbb{E}(\cdot)$ indicates the statistical expectation. The symbol $\otimes$ is used to represent the Kronecker product. Lowercase bold letters (e.g., $\mathbf{y}$) represent vectors, uppercase bold letters (e.g., $\mathbf{Y}$) represent matrices, and non-bold letters denote scalars.
@@ II System Model
We consider a real-world scenario where a signal propagates through a three-dimensional (3D) space, with a local spherical coordinate system defined such that $\phi$ represents the azimuth angle and $\psi$ denotes the elevation angle [sp1]. The reflecting and antenna elements are arranged in a two-dimensional grid within the YZ plane. Each element has $d_{H}$ and $d_{V}$ corresponding to the horizontal and vertical spacing, respectively. To systematically index the elements, they are ordered row by row and denoted as $n\in[1,N]$ RIS elements and $m\in[1,M]$ antenna elements. The $n$-th element’s position relative to the origin is:
@@
where $\text{mod}(\cdot,\cdot)$ denotes the modulus operation, and $\lfloor\cdot\rfloor$ represents the floor function. Now, consider a plane wave with wavelength $\lambda$ impinging on the RIS from an azimuth angle $\phi$ and an elevation angle $\psi$. The array response vector corresponding to this incident wave is given by:
@@
$\boldsymbol{\omega}(\phi,\psi)=\frac{2\pi}{\lambda}\begin{bmatrix}\cos(\psi)\cos(\phi)\\ \cos(\psi)\sin(\phi)\\ \sin(\psi)\end{bmatrix}$. (4)
@@
$\mathbf{f}_{k}=\sqrt{\frac{\eta_{r}}{\eta_{r}+1}}\mathbf{f}^{{LoS}}_{k}+\sqrt{\frac{1}{\eta_{r}+1}}\mathbf{f}^{{NLoS}}_{k}.$ (5)
@@
where $\alpha_{n}$ and $\gamma_{n}$ are the indices of reflecting elements in the horizontal and vertical directions, respectively. We consider the Non-LoS (NLoS) channels as quasi-static Rayleigh fading, which remains constant for the coherence interval. The NLoS channel between the RIS and the user is modeled using correlated Rayleigh fading [sp0]. The NLoS channel from the RIS to the user is given by:
@@ II-A Channel Estimation Framework
The system’s performance is heavily dependent on the availability of precise channel information. We adopted a Time Division Duplex (TDD) protocol, which exploits channel reciprocity to estimate downlink CSI from uplink CSI. The TDD protocol divides communication into time slots, known as coherence intervals, during which the channel response remains approximately constant [sp0]. We divide the channel estimation framework into two phases: direct channel estimation and cascaded channel estimation. In the first phase, the RIS is turned off, so no reflections occur. The direct channel between the user and the BS is estimated during this stage, as shown in Fig. 2. Moreover, in the second phase, we estimate the cascaded channel between the user, the RIS, and the BS. Estimating the cascaded channel requires $L$ distinct phase shift matrices, where $L\geq N$, to facilitate accurate channel estimation. For each phase shift matrix, each user is assigned an orthogonal pilot sequence $\mathbf{x}_{k}$ of length $u$, where $u\geq K$, ensuring the pilot’s orthogonality for accurate estimation. For the $k$-th user during the $l$-th subframe, the pilot signal received at the BS, denoted as $\mathbf{y}_{k}\in\mathbb{C}^{M\times 1}$, is given by the following expression:
@@
$\mathbf{y}_{k}=\frac{1}{Pu}(\mathbf{b}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{H}\text{diag}(\mathbf{\Theta}_{l})\mathbf{f}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{n}_{k}\mathbf{x}_{k}).$ (15)
@@
$\mathbf{y}_{k}=\mathbf{b}_{k}+\mathbf{H}\text{\text{diag}}(\mathbf{\Theta}_{l})\mathbf{f}_{k}+\mathbf{v}_{k}$ (16)
@@
where $\mathbf{G}_{k}=\ \mathbf{H}\,\text{diag}(\mathbf{f}_{k})\in\mathcal{C}^{M\times N}$ represents the cascaded channel and $\mathbf{V}_{k}=[\mathbf{v}_{k,1},\mathbf{v}_{k,2},\ldots,\mathbf{v}_{k,L}]$. According to [DL4], designing the optimal phase shift matrix $\mathbf{S}$ using a discrete Fourier transform maximizes the received signal power at the BS, which can be expressed as follows:
@@
where $W=e^{j\frac{2\pi}{L}}$ and $\mathbf{S}\mathbf{S}^{H}=L\mathbf{I}_{N}$. Cascaded channel estimation involves recovering $\mathbf{G_{k}}$ from the received pilots. The simplest approach is the LS method [LS], which performs effectively when the channel is assumed to be static. The LS representation of the cascaded channel can be expressed as $\mathbf{\tilde{G}}_{k}^{LS}=\mathbf{Y}_{k}\mathbf{S}^{H}(\mathbf{S}\mathbf{S}^{H})^{-1}$, where $\mathbf{\tilde{G}}^{LS}=[\mathbf{\tilde{G}}_{1}^{LS},\mathbf{\tilde{G}}_{2}^{LS},\ldots,\mathbf{\tilde{G}}_{\tau}^{LS}]$, and where $\tau$ represent total training data. However, treating the channel as static fails to capture its spatial characteristics, especially when compared to data-driven methods. Moreover, under a Rayleigh fading model and assuming the availability of the statistical channel correlation matrix, the optimal estimator is the MMSE estimator. However, when the channel distribution is complex, as in the case of the cascaded channel in our scenario, the MMSE estimator becomes computationally expensive. To reduce this complexity, the MMSE estimator is often approximated by the LMMSE estimator. According to [LMMSE], the LMMSE estimator is given by:
@@
where $\mathbf{R}_{G_{k}}=\mathbb{E}(G_{k}^{H}G_{k})$ denotes the statistical channel correlation matrix. However, since $G_{k}$ corresponds to the kthk-th user’s cascaded channel, it does not follow the Rayleigh fading model. Consequently, the assumptions underlying the LMMSE estimator are violated, leading to a significant performance gap between the LMMSE estimation and the actual channel. To address this limitation, we propose a deep learning-based method that effectively learns the underlying channel structure and mitigates the performance gap introduced by the model mismatch in conventional methods.
@@ III Deep Learning Assisted Channel Estimation
Deep learning-assisted cascaded channel estimation has gained significant popularity in recent years [DL5, DL1, DL2, DL3, DL4, DL6]. One of the key reasons for this growing interest is its extremely low inference time, which makes it well-suited for low-latency applications. However, existing works primarily focus on small-scale scenarios with only a limited number of antennas and RIS elements. In this paper, we investigate the use of deep learning to estimate the channel in RIS-enabled XL-MIMO systems, which may involve hundreds of RIS elements and thousands of antennas. We leverage the spatial correlation inherent in the cascaded channel, resulting from the closely placed RIS elements and BS antennas. Building on this property, we propose a patch-based training framework where small, localized patches are extracted from each high-dimensional data point $\mathbf{\tilde{G}}^{\mathrm{LS}}_{i}\in\mathbb{C}^{M\times N}$, capturing salient features while reducing input dimensionality and enabling scalable training for large-scale systems on resource-constrained edge devices. This section presents the proposed deep learning framework, while Sections IV and V discuss spatial correlation and the patch-based training strategy.
@@ III-A Theoretical Background
Recent studies have demonstrated that encoder–decoder architectures can effectively learn compact, low-dimensional representations by capturing essential patterns and structures while minimizing reconstruction loss [AE1, UNET1, GM2]. These architectures also exhibit strong generalization capabilities, making them suitable for modeling complex, high-dimensional data distributions. Moreover, they typically require less training data, as they learn latent representations that preserve the most critical features of the input. This latent space not only facilitates efficient data compression and generalization but also enables the extraction of meaningful features, which is particularly beneficial in our case. Despite their effectiveness, encoder–decoder models may face reconstruction challenges, as compression into the bottleneck layer can result in the loss of crucial spatial features, limiting the recovery of fine-grained structures.
To address this issue, our proposed architecture incorporates skip connections that directly transfer high-resolution features from the encoder to the decoder [UNET1]. This design preserves essential spatial correlations while simultaneously filtering out noise [GM2]. In addition to skip connections, our framework further enhances the architecture by aggregating feature maps from multiple hierarchical levels, thereby improving upon the standard model through multi-scale feature integration. This approach enables the network to capture richer contextual information, effectively combining fine-grained local details with broader structural patterns. The enhanced feature aggregation substantially improves estimation performance by preserving critical spatial structures and suppressing noise more effectively, even in scenarios with limited training data. Consequently, the proposed architecture offers a more robust and efficient solution for denoising tasks in large-scale systems.
Figure 3: The proposed framework employs $l$ levels, each with a dedicated denoising block in the encoder. Filters double, and the resolution halves at each level. Fused encoder and decoder features serve as input to each decoder layer.
@@ III-B Proposed Model Architecture
Our proposed architecture is designed to effectively capture the underlying signal features present in the dataset. The framework integrates a dedicated denoising module that employs a subtraction-based structure to enhance channel estimation by isolating the core signal components from additive noise. The architecture, illustrated in Fig. 3, adopts a multi-level encoder–decoder design to exploit spatial hierarchies to break the noise correlation, enabling more effective denoising and improved estimation accuracy. In the encoder, each level in the architecture consists of multi-layer Convolutional Neural Networks (CNNs) with distinct denoising blocks. The output of each higher-level block is downsampled by a factor of two and fed into the subsequent lower-level block, where the number of filters is doubled to capture increasingly complex features. We employ a MaxPooling layer with a (2,2) kernel in the encoder to downsample between levels. In the decoder, the output from lower-level blocks is upsampled to the current level, whereas the feature maps from the corresponding encoder blocks are downsampled to match that level. Additionally, the output from the same level in the encoder is concatenated with these other outputs, enabling effective multi-scale information fusion to preserve spatial structure. Moreover, we utilized bilinear interpolation for upsampling between levels in the network to smoothly increase the resolution of feature maps while maintaining spatial consistency. During the decoder phase, the filter size is halved, which helps to refine the reconstructed output. Finally, an activation function is applied to the aggregated features, adding nonlinearity and improving the model’s learning of complex mappings before the final output.
@@ III-C Denoising Block Structure
The CNN architecture is well-suited to extract underlying signal features from noisy data effectively. In [DL1], a subtraction-based mechanism was proposed to eliminate noise from the data, demonstrating effectiveness in denoising. However, this approach was restricted to learning models with identical dimensions, which is incompatible with the design of our proposed architecture. To address this limitation, the proposed denoising architecture introduces a $1\times 1$ convolutional layer at the beginning of each level in the encoder to align the input and output dimensions, thereby facilitating the application of the denoising mechanism. The denoising block architecture is composed of a CNN with $i$ layers, where each layer consists of a convolution operation, batch normalization, and a non-linear activation function, as illustrated in Fig. 3. The denoising block applies a pointwise ($1\times 1$) convolution to align the input with the network dimensions, followed by multiple convolutional layers. Finally, an element-wise subtraction is performed between the $l$-th level’s block input $\mathbf{I}_{l}$ and output $\tilde{\mathbf{z}}_{\theta}(\mathbf{I}_{l})$ to produce the denoised channel matrix, effectively preserving essential features. At the $l$-th level, it can be expressed as:
$\mathbf{Z}_{l}=\mathbf{I}_{l}-\tilde{\mathbf{Z}}_{\theta}(\mathbf{I}_{l}),$ (20)
where $\tilde{\mathbf{Z}}_{\theta}$ denotes the convolutional neural network and $\theta$ represents its parameters. The primary challenge associated with cascaded channel estimation is the presence of Additive White Gaussian Noise (AWGN) [OP1]. The simplest estimator for this task is the LS method, as it does not require statistical information. However, channel estimation using LS is often inadequate, as it is highly sensitive to noise, particularly in low Signal-to-Noise Ratio (SNR) conditions. Although it may still provide a rough estimate, its performance significantly degrades in noisy environments. Therefore, we use the LS estimator as input to our model, as it allows for a coarse estimate of the cascaded channel [DL1]. This initial estimate enables our model to focus on refining the channel representation, rather than learning it from scratch, thereby improving convergence and overall performance. Moreover, to accommodate the complex-valued nature of the data, each data point is decomposed into its real and imaginary components and provided as separate input channels to facilitate effective feature extraction.